# url]https://orcid.org/$0$-0001-7408-6925

### 26 Nov 2023 v1

@@ Abstract
The application of Machine learning to finance has become a familiar approach, even more so in stock market forecasting. The stock market is highly volatile and huge amounts of data are generated every minute globally. The extraction of effective intelligence from this data is of critical importance. However, a collaboration of numerical stock data with qualitative text data can be a challenging task. In this work, we accomplish this and provide an unprecedented, publicly available dataset with technical and fundamental data, sentiment that we gathered from News Archives, TV news captions, Radio Transcripts, Tweets, Daily financial newspapers, etc. The text data entries used for sentiment extraction total more than 1.4 Million. The dataset comprises of daily entries from January 2018 to December 2022 for 8 different companies and Dow Jones Index as a whole. Holistic Fundamental and Technical data is provided training ready for Model learning and deployment. The predictive power of deep learning models is highly determined by the training data provided. This dataset would be of benefit for research globally incorporating qualitative intelligence for stock market forecasting.The dataset is made available at Github.
@@
A robust dataset, with quantitative and qualitative data for stock forecasting.
@@ 1 Introduction
However, there are challenges to extracting information from sparse web data. First, getting benchmark, recognized data is difficult. Even though the availability of text data is soaring, it is usually limited to prominent stocks and might not be of a good standard. Additionally, data is sparse meaning getting daily sentiments for stocks can prove to be difficult. Further, extraction has to be performed from unstructured web sources which has restricted by IP blocking. Next, both events and emotions contribute to a qualitative understanding of financial markets. Blooming progress in NLP has given researchers a new purview on stock market prediction by extraction of investor emotions and moods[7][8].
It is arduous and impractical for investors or researchers to analyze big data for predictions. Added to this there is a heavy requirement of domain knowledge in finance for the selection of data parameters. The purpose of the paper is to provide a publicly available data set for stock market prediction using qualitative data that is ready for ML model learning that could be implemented by the layman Machine Learning research community. The dataset comprises of 8 stocks from diverse industries. Added to this, Dow Jones Industrial Average(DJIA) data is also provided. In our dataset, we make use of financial articles from recognized sources online to mitigate fake news. The main contributions of the study to the research community are summarised as follows:
@@
    Substantial data collection sources and techniques including API’s, twitter data, radio transcripts, and news archive data.
@@ 2 Related Work
Early proceedings debated the possibility of prediction of the stock market. According to the Random Walk theory, stock prices move stochastically in a random walk, making it impossible to forecast and indicating that any attempt at prediction would be unsuccessful. [9]. Next, the Efficient market hypothesis states that the market is random but the market is also efficient in determining the optimal stock market price. However, further research in the field has shown that stock price direction analysis is possible.[10][11] However, as raised before historical stock data has limitations in unfolding all the information relating to firms’ financial well-being.
@@
In this study forecasting and predictions have been used interchangeably. Stock market prediction was also based on different time frames. These time frames can be established for intraday or inter-day trading. Most research aimed at making forecasting for stock movement for one next day [12][14][15]. Gong et al. in [20] aimed at making monthly directional predictions. Patel et al. in [19] made predictions on a range of time frames from 1 to 30 days. Nassirtoussi et al. in [18] implemented predictions every 20 minutes.
As aforementioned, the analysis of stocks happens through two categories, Qualitative and Quantitative. Quantitative analysis includes two sub-categories: Technical indicators and Quantitative Fundamental Analysis. Quantitative analysis typically consists of Qualitative Fundamental analysis. The following section elucidates some current and pertinent literature based on each category.
@@ 2.1.1 Technical Indicators
Technical analysis of stocks aims at understanding stock behaviour by analyzing statistical trends, patterns, and historical market data. It is founded on the idea that a financial time series’ future behaviour is conditioned by its own history. In the wide range of research performed on Qualitative data, different studies make use of different technical indicators[21], partly because of different indexes used. Some studies used common technical indicators in stock markets [22], while others used a more diverse range of indicators[23]. In this work, we perform an extensive review of which technical indicators are more effective in relation to different indexes and comprehend the best indicators to be included in our Dataset. The review is demonstrated in Table(9).
@@ 2.1.2 Quantitative Fundamental analysis
Fundamental analysis of stocks is a method used by investors to evaluate the intrinsic value of a company by examining its underlying financial and qualitative factors. Quantitative Fundamental analysis is based on the usage of numerical data and Qualitative Fundamental analysis makes use of text data. Technical analysis tactics simply take into account a stock’s price history; Quantitative fundamental analysis techniques take into account a company’s financial condition as well as macroeconomic variables [24]. Even though advocates of EMH contend that a stock’s intrinsic value is always equal to its present price, fundamental analysts purchase and sell companies when the intrinsic value is lower or higher than the market price.
@@
Studies have concluded that regional macroeconomic factors also have a significant impact on how the equity market performs. According to Zhao et al [25], the trajectory of the A-share markets in mainland China can be impacted by changes in the RMB exchange rate. Consequently, including macroeconomic variables will provide more intelligence for neural network training for better training.
@@
Typically, fundamental research is conducted from “a macro to micro viewpoint” to pinpoint securities/assets that the marketplace has not accurately valued. Analysts often research the following topics in the following order:
@@
In our study, we made use of Macroeconomic factors, (shown under Green in Table(10)) and Financial Ratios. Financial Ratios of stocks refer to the quantitative metrics or factors associated with a company’s financial performance and position. All the different ratios and Macroeconomic factors are explained in Table(10) with colour coding for readability. Different Financial Ratios can be categorized into([27]):
@@ 2.2 Qualitative Data-Based Studies
Qualitative data analysis in the context of stocks involves examining non-numerical information about companies or the market to make investment decisions. This typically involves analysis of text data such as news articles, tweets, financial forum discussions etc. Multiple studies have shown that The stock price can be significantly impacted by financial news [28][29]. Some text mining methods have been summarized in [30]. Investigations have been performed, utilizing sentiment analysis to assess the influence of sentiments on stock market volatility. Also, it was shown that the Volatility Index(VIX) is highly correlated with market sentiment[31]. Preceding research analyzed the role of Author tone from financial articles and evaluated its sentiment[32]. A combined effect of web text news articles and social media sentiment was taken into account in [33].
Social media has been used extensively by numerous research works through time.[34][35]. Tools such as Tweety, OpinionFinder, etc can be used for the extraction and analysis of Twitter data. It’s proven already that Twitter’s mood impacts stock market prediction[15]. In [36], sentiment and survey indices( USMC, AAII, and Sentix) were extracted from large Twitter data for the S&P500 index. To integrate survey sources and blogs, the Kalman filter was utilized. Finally, for metrics, Diebold-Mariano test was performed to check for a significant difference.
Initially, some studies used the bag of words approach for embedding news text data for prediction, but Xiao Ding in [37] showed that, Unlike the bag of words approach, structured events are more helpful depictions for the purpose of making stock market predictions. He also concluded that in the study of forecasting, the quality of the information is more significant than the amount of data available. Moreover, studies have investigated the fusion of event and sentiment mood data for prediction. In [38] sentiments and events are integrated and exhibit combined influence on the stock market making use of tensor framework.
A recent work by Zhang .et al [1] effective integration of events and mood data was implemented addressing the sparse data problem. Studies in Behavioural Finance that correlate with investor decision-making processes have discussed the implications of overconfidence [39]. According to earlier studies on overconfidence and behavioural psychology [40] it was stated that investors choose private knowledge over public [41]. Thus because of investors’ access to social media information, overconfidence may be reduced. Added to this, a comprehensive review of Qualitative Data-Based Studies is shown in Table(12).
@@ 3 Dataset
In this section, we propose a Huge stock market forecasting dataset consisting of features in three categories; Technical Indicators, qualitative fundamental indicators, and Quantitative fundamental indicators. We consider 8 different stocks shown in Table(1) from DJIA (Dow Jones Industrial Average) and provide data for them. Data is also provided for the de facto DJIA index. It is to be noted that Dow Jones is a price-weighted Index.
@@ 3.2 Quantitative Fundamental Analysis Features
We added Financial Variables and other Macroeconomic Variables from 10 using various sources for each stock under consideration. The purpose of Quantitative fundamental analysis as mentioned earlier is to assess the financial health of these stocks through Financial firm-specific data. Data comprises of company Income statements, Balance Sheets, Cash flows, Commodities, Macro variables, multiple financial ratios, International currency exchange rates, and other indexes. In total, we extracted around 220 features including technical and Fundamental Qualitative analysis for each stock. Since Dow Jones data does not include company-specific fundamental data hence consists of around 120 features without the inclusion of Qualitative Analysis.
@@ 3.3.1 Data Collection Module
In the world of big data, there is a large amount of data available on the web. Extraction can be performed through various techniques. There are pre-designed API’s for this task. However, APIs can only be used to extract data through queries to a specific data provider/website. So in our work, we design a personalized system for data extraction from multiple sources. We develop a Web Scraper for data collection from Internet Archive. This provided us with quality data per day from January 2018 to December 2022 for each year. The data was stock-specific to all 8 stocks and lastly "Dow Jones" specific. The data collected was of two types, TV news captions data and Radio Transcripts data. Web scraping was hedged against IP blocking utilizing IP rerouting and random interval multi-threading using Redis-DB. The respective search engine is made open source at Github.
@@
Next, we used the New York Times financial newspaper API which provides official news articles. The reason for choosing this API is essentially because NYT provides free access and the API provides date-annotated news articles. Data can also be extracted from various other web news providers like the Financial Times, Thomas Reuters, Refinitiv, and Washington Post. This provided us with pre-structured data for daily news in the same time frame. However here we encountered some of the biggest problems with the research which was to find APIs with the capability to query news articles in a specific time frame for a specific stock in the category of Business or Fianance. Information that does lead to intelligence is useless. This would also confuse machine learning models trained on our data after sentiment analysis was performed on news articles unrelated to a specific stock. The problem with the New York Times was that even though we could extract about 2k large articles, the data wasn’t recognized for a specific stock. To work around this we try keyword matching shown in 3.3.2 but with failure. Hence we made use of 3 other APIs; Guardian API, Current API, and Alpha Vantage news API. We manually tweaked the parameters of the API for the best retrieval of most related responses in terms of stock entity similarity. Even though these APIs didn’t contribute to consistent data(was sparse) through the 5 years they do provide us with valuable news articles in a more scattered fashion.
Finally, we utilize Twitter data. A self-designed Twitter API based on Selenium and SQLite3 was utilized for date and keyword-specific data extraction. The purpose of the preparation of the code was to utilize as many attributes as possible and get intelligent task-specific data instead of inexact data. This solved the problem of time-specific and stock-specific information retrieval. The code also works with multiple threads and hedges against IP blocking. In [42] it was concluded that Twitter attributes such as likes, followers, retweet number, and account verification add to better predictability of models. Hence we also capture likes, comment numbers, and retweet numbers along with the post date and Twitter body. We used each stock name and Dow Jones as a query. We capture 0.83 Million Tweets in total. Twitter extraction engine is available at Github.
@@ 3.3.2 Keyword matching
The structure of the response from NYT API is as follows: Snippet, Web URL, Lead Paragraph, Headline, Keywords, and variables such as type of material and publish date. We extracted 45773 articles in 5 years. Our focus will be on keywords, NYT already provides us with extracted keywords from each news article, however, there is no generic standard form for these words. Hence to co-relate the news articles with each ticker from the ones selected in 1 we have to match them. For this, we utilize cosine similarity and edit distance, an approach taken in [43]. After capturing different keywords for each ticker, we then extract stock-specific data per ticker. To capture articles related to Dow Jones directly, we took keywords related to "Dow Jones". The results were very sparse with only 21 articles from 45773 being extracted. The result for each stock index were sparse too as expected[43]. We could only get 212 articles that were related to any of the eight stocks from 5 years. Thus, keyword matching proved to be insufficient.
@@ 3.3.3 Preprocessing
For Quantitative data, we don’t apply any preprocessing because we want to make the dataset as flexible as possible so that research would be able to apply any type and number of techniques on the original untreated dataset.
However, for Qualitative data, it is vital that optimal preprocessing of text is performed. This is even more crucial with twitter data where informal lingo is used by the users. Misspelled words, slang, New words, URLs, etc are common [44]. Hence, additional preprocessing has to be applied for Twitter data over news and archive data. Further, stemming is to be performed for all news and archive data when used with Loughran-McDonald Master Dictionary Sentiment Analysis. Loughran-McDonald method is elucidated below. This is because all different forms of words might not be present in the dictionary, so it’s optimal all the words are converted to root form. This way when direct String comparisons are done the words are matched with the dictionary appropriately. For example, the word "boom" is present in the dictionary, "booming" is not and unless stemming is used this conversion does not happen. The pipeline used for preprocessing is shown in Figure3. It is to be noted that we don’t remove "@organization" words such as "@Apple" but only remove the "@" token from the corpus. Corpus is also not converted to lowercase, this is to prevent the loss of user intonation and speech intensity information in text.

### 01 Jan 2024 v3

@@ Abstract
The application of Machine learning to finance has become a familiar approach, even more so in stock market forecasting. The stock market is highly volatile, and huge amounts of data are generated every minute globally. The extraction of effective intelligence from this data is of critical importance. However, a collaboration of numerical stock data with qualitative text data can be a challenging task. In this work, we accomplish this by providing an unprecedented, publicly available dataset with technical and fundamental data and sentiment that we gathered from news archives, TV news captions, radio transcripts, tweets, daily financial newspapers, etc. The text data entries used for sentiment extraction total more than 1.4 Million. The dataset consists of daily entries from January 2018 to December 2022 for eight companies representing diverse industrial sectors and the Dow Jones Industrial Average (DJIA) as a whole. Holistic Fundamental and Technical data is provided training ready for Model learning and deployment. Most importantly, the data generated could be used for incremental online learning with real-time data points retrieved daily since no stagnant data was utilized. All the data was retired from APIs or self-designed robust information retrieval technologies. These adaptable technologies facilitate data extraction for any stock. Moreover, the utilization of Spearman’s rank correlation over real-time data, linking stock returns with sentiment analysis has produced noteworthy results for the DJIA, achieving accuracy levels surpassing 60%. The dataset is made available at Github.
@@
A robust dataset, with quantitative and qualitative data for Incremental Online learning at stock forecasting.
@@ 1 Introduction
However, there are challenges to extracting information from sparse web data. First, getting benchmark, recognized data is difficult. Even though the availability of text data is soaring, it is usually limited to prominent stocks and might not be of a good standard. Additionally, data is sparse meaning getting daily sentiments for stocks can prove to be difficult. Further, extraction has to be performed from unstructured web sources which are restricted by IP blocking. Next, both events and emotions contribute to a qualitative understanding of financial markets. Blooming progress in NLP has given researchers a new purview on stock market prediction by extraction of investor emotions and moods[7][8].
It is arduous and impractical for investors or researchers to analyze big data for predictions. Added to this there is a heavy requirement of domain knowledge in finance for the selection of data parameters. The purpose of the paper is to provide a publicly available data set for stock market prediction using qualitative data that is ready for ML model learning that could be implemented by the layman Machine Learning research community. The dataset comprises 8 stocks from diverse industries. Added to this, Dow Jones Industrial Average(DJIA) data is also provided. In our dataset, we make use of financial articles from recognized sources online to mitigate fake news. The main contributions of the study to the research community are summarised as follows:
@@
    Substantial data collection sources and techniques including APIs, twitter data, radio transcripts, and news archive data.
@@ 2 Related Work
Early proceedings debated the possibility of predicting the stock market. According to the Random Walk theory, stock prices move stochastically in a random walk, making it impossible to forecast and indicating that any attempt at prediction would be unsuccessful. [9]. Next, the efficient market hypothesis states that the market is random but also efficient in determining the optimal stock market price. However, further research in the field has shown that stock price direction analysis is possible.[10][11] However, as raised before historical stock data has limitations in unfolding all the information relating to firms’ financial well-being.
@@
In this study, forecasting and predictions are used interchangeably. Stock market prediction was also based on different time frames. These time frames can be established for intraday or inter-day trading. Most research aimed at making forecasting for stock movement for one next day [12][14][15]. Gong et al. in [20] aimed at making monthly directional predictions. Patel et al. in [19] made predictions on a range of time frames from 1 to 30 days. Nassirtoussi et al. in [18] implemented predictions every 20 minutes.
As mentioned, the analysis of stocks happens through two categories, qualitative and quantitative. Quantitative analysis includes two sub-categories: Technical indicators and Quantitative Fundamental Analysis. Quantitative analysis typically consists of Qualitative Fundamental analysis. The following section elucidates some current and pertinent literature based on each category.
@@ 2.1.1 Technical Indicators
Technical analysis of stocks aims to understand stock behavior by analyzing statistical trends, patterns, and historical market data. It is founded on the idea that a financial time series’ future behavior is conditioned by its own history. In the wide range of research performed on Qualitative data, different studies make use of different technical indicators[21], partly because of different indexes used. Some studies used common technical indicators in stock markets [22], while others used a more diverse range of indicators[23]. In this work, we perform an extensive review of which technical indicators are more effective in relation to different indexes and comprehend the best indicators to be included in our Dataset. The review is demonstrated in Table(9).
@@ 2.1.2 Quantitative Fundamental analysis
Fundamental analysis of stocks is a method used by investors to evaluate the intrinsic value of a company by examining its underlying financial and qualitative factors. Quantitative Fundamental analysis is based on the usage of numerical data, and Qualitative Fundamental analysis makes use of text data. Technical analysis tactics simply take into account a stock’s price history; Quantitative fundamental analysis techniques take into account a company’s financial condition as well as macroeconomic variables [24]. Even though advocates of EMH contend that a stock’s intrinsic value is always equal to its present price, fundamental analysts purchase and sell companies when the intrinsic value is lower or higher than the market price.
@@
Studies have concluded that regional macroeconomic factors also have a significant impact on how the equity market performs. According to Zhao et al. [25], the trajectory of the A-share markets in mainland China can be impacted by changes in the RMB exchange rate. Consequently, including macroeconomic variables will provide more intelligence for neural network training for better training.
@@
Typically, fundamental research is conducted from “a macro to the micro viewpoint” to pinpoint securities/assets that the marketplace has not accurately valued. Analysts often research the following topics in the following order:
@@
In our study, we made use of Macroeconomic factors, (shown under Green in Table(10)) and Financial Ratios. Financial Ratios of stocks refer to the quantitative metrics or factors associated with a company’s financial performance and position. All the different ratios and Macroeconomic factors are explained in Table(10) with color coding for readability. Different Financial Ratios can be categorized into([27]):
@@ 2.2 Qualitative Data-Based Studies
Qualitative data analysis in the context of stocks involves examining non-numerical information about companies or the market to make investment decisions. This typically involves analysis of text data such as news articles, tweets, financial forum discussions, etc. Multiple studies have shown that The stock price can be significantly impacted by financial news [28][29]. Some text mining methods have been summarized in [30]. Investigations have been performed, utilizing sentiment analysis to assess the influence of sentiments on stock market volatility. Also, it was shown that the Volatility Index(VIX) is highly correlated with market sentiment[31]. Preceding research analyzed the role of Author tone from financial articles and evaluated its sentiment[32]. A combined effect of web text news articles and social media sentiment was taken into account in [33].
Social media has been used extensively by numerous research works through time.[34][35]. Tools such as Tweety, OpinionFinder, etc can be used for the extraction and analysis of Twitter data. It’s proven already that Twitter’s mood impacts stock market prediction[15]. In [36], sentiment and survey indices( USMC, AAII, and Sentix) were extracted from large Twitter data for the S&P500 index. The Kalman filter was utilized to integrate survey sources and blogs. Finally, for metrics, Diebold-Mariano test was performed to check for a significant difference.
Initially, some studies used the bag of words approach for embedding news text data for prediction, but Xiao Ding in [37] showed that, Unlike the bag of words approach, structured events are more helpful depictions for the purpose of making stock market predictions. He also concluded that in the study of forecasting, the quality of the information is more significant than the amount of data available. Moreover, studies have investigated the fusion of event and sentiment mood data for prediction. In [38] sentiments and events are integrated and exhibit combined influence on the stock market, making use of the tensor framework.
A recent work by Zhang .et al [1] effective integration of events and mood data was implemented addressing the sparse data problem. Studies in Behavioural Finance that correlate with investor decision-making processes have discussed the implications of overconfidence [39]. According to earlier studies on overconfidence and behavioral psychology [40] it was stated that investors choose private knowledge over public [41]. Thus, because of investors’ access to social media information, overconfidence may be reduced. Added to this, a comprehensive review of Qualitative Data-Based Studies is shown in Table(12).
@@ 3 Dataset
In this section, we propose a Huge stock market forecasting dataset consisting of features in three categories: Technical Indicators, qualitative fundamental indicators, and Quantitative fundamental indicators. We consider eight different stocks shown in Table(1) from DJIA (Dow Jones Industrial Average) and provide data for them. Data is also provided for the de facto DJIA index. It is to be noted that Dow Jones is a price-weighted Index.
@@ 3.2 Quantitative Fundamental Analysis Features
We added Financial Variables and other Macroeconomic Variables from 10 using various sources for each stock under consideration. The purpose of Quantitative fundamental analysis as mentioned earlier is to assess the financial health of these stocks through Financial firm-specific data. Data comprises company Income statements, Balance Sheets, Cash flows, Commodities, Macro variables, multiple financial ratios, International currency exchange rates, and other indexes. In total, we extracted around 220 features, including technical and fundamental qualitative analyses for each stock. Since Dow Jones data does not include company-specific fundamental data, it consists of around 120 features without the inclusion of Qualitative Analysis.
@@ 3.3.1 Data Collection Module
In the world of big data, there is a large amount of data available on the web. Extraction can be performed through various techniques. There are pre-designed API’s for this task. However, APIs can only be used to extract data through queries to a specific data provider/website. So, in our work, we design a personalized system for data extraction from multiple sources. We develop a Web Scraper for data collection from Internet Archive. This provided us with quality data per day from January 2018 to December 2022 for each year. The data was stock-specific to all eight stocks and lastly "Dow Jones" specific. The data collected was of two types: TV news captions data and Radio Transcripts data. Web scraping was hedged against IP blocking utilizing IP rerouting and random interval multi-threading using Redis-DB. The respective search engine is made open source at Github.
@@
Next, we used the New York Times financial newspaper API, which provides official news articles. The reason for choosing, this API is essentially because NYT provides free access and the API provides date-annotated news articles. Data can also be extracted from various other web news providers like the Financial Times, Thomas Reuters, Refinitiv, and Washington Post. This provided us with pre-structured data for daily news in the same time frame. However here we encountered some of the most challenging problems with the research which was to find APIs with the capability to query news articles in a specific time frame for a specific stock in the category of Business or Fianance. Information that does lead to intelligence is useless. This would also confuse machine learning models trained on our data after sentiment analysis was performed on news articles unrelated to a specific stock. The problem with the New York Times was that even though we could extract about 2k large articles, the data wasn’t recognized for a specific stock. To work around this we try keyword matching shown in 3.3.2 but with failure. Hence, we made use of 3 other APIs: Guardian API, Current API, and Alpha Vantage news API. We manually tweaked the parameters of the API for the best retrieval of most related responses in terms of stock entity similarity. Even though these APIs didn’t contribute to consistent data(was sparse) through the five years they do provide us with valuable news articles in a more scattered fashion. Next, IP blockage was a challenge for web scraping for radio transcripts and news archive data. To work around this, we developed robust, privacy-aware retrieval methods that leverage IP rotation via proxy servers, random time interval scheduling, and user-agent switching using Redis cache to replicate human behavior.
Finally, we utilize Twitter data. A self-designed Twitter API based on Selenium and SQLite3 was utilized for date and keyword-specific data extraction. The purpose of the code preparation was to utilize as many attributes as possible and get intelligent task-specific data instead of inexact data. This solved the problem of time-specific and stock-specific information retrieval. The code also works with multiple threads and hedges against IP blocking. In [42] it was concluded that Twitter attributes such as likes, followers, retweet number, and account verification add to better predictability of models. Hence, we also capture likes, comment numbers, and retweet numbers along with the post date and Twitter body. We used each stock name and Dow Jones as a query. We capture 0.83 Million Tweets in total. Twitter extraction engine is available at Github.
@@ 3.3.2 Keyword matching
The structure of the response from NYT API is as follows: Snippet, Web URL, Lead Paragraph, Headline, Keywords, and variables such as type of material and publish date. We extracted 45773 articles in 5 years. Our focus will be on keywords, NYT already provides us with extracted keywords from each news article, however, there is no generic standard form for these words. Hence, to co-relate the news articles with each ticker from the ones selected in 1 we have to match them. For this, we utilize cosine similarity and edit distance, an approach taken in [43]. After capturing different keywords for each ticker, we then extract stock-specific data per ticker. To capture articles related to Dow Jones directly, we took keywords related to "Dow Jones". The results were very sparse with only 21 articles from 45773 being extracted. The result for each stock index were sparse too as expected[43]. We could only get 212 articles that were related to any of the eight stocks from 5 years. Thus, keyword matching proved to be insufficient.
@@ 3.3.3 Preprocessing
For quantitative data, we don’t apply any preprocessing because we want to make the dataset as flexible as possible so that research can apply any type and number of techniques to the original untreated dataset.
However, for Qualitative data, it is vital that optimal preprocessing of text is performed. This is even more crucial with Twitter data where informal lingo is used by the users. Misspelled words, slang, New words, URLs, etc are common [44]. Hence, additional preprocessing has to be applied for Twitter data over news and archive data. Further, stemming is to be performed for all news and archive data when used with Loughran-McDonald Master Dictionary Sentiment Analysis. Loughran-McDonald method is elucidated below. This is because all different forms of words might not be present in the dictionary, so it’s optimal for all the words to be converted to root form. This way when direct String comparisons are done the words are matched with the dictionary appropriately. For example, the word "boom" is present in the dictionary, but "booming" is not, and unless stemming is used this conversion does not happen. The pipeline used for preprocessing is shown in Figure3. It is to be noted that we don’t remove "@organization" words such as "@Apple" but only remove the "@" token from the corpus. Corpus is also not converted to lowercase, this is to prevent the loss of user intonation and speech intensity information in text.